1. a coin is tossed 3 times. find the probability of obtaining
A.) heads on the first and last toss, and tails on the second toss.
B.) at least 2 heads
C.) at most 2 heads
2. show that the median of the set of numbers 1, 2, 4, 8, 16, 32 is 6. how does this compare with the mean?
3. a jar contains 7 red, 6 green, 8 blue, and 4 yellow marbles. a marble is chosen at random from the jar. after replacing it, a second marble is chosen. what is the probability of choosing: A.) 2 yellow marbles? B.) at least one red marble?

Answer by Boreal(15235)

1-3-3-1: 0 heads (1), 1 head (3) 2 heads (3), 3 heads (1). At least 2 heads is 4/8 or 1/2.
HHH
HHT
HTH
HTT
THH
THT
TTH
TTT
HTH has a probability of 1/8
At least two heads is 1/2
At most 2 heads is 7/8.
=======================
1==2==4==8==16==32
The median is the middle of 6, which has no defined middle term. Therefore, it is the average of the terms on either side of the middle, which are 4 and 8. Their average is 6. The mean is the sum divided by 6, or 10.5. This is higher, because the distribution has two numbers (16 and 32) significantly larger than the others.
======================
25 marbles altogether
2 yellow is 4/25*4/25=16/625 or 0.0256
at least one red marble is 7/25*18/25+18/25*7/25+7/25*7/25, which is probability of both being red.
that is (126+126+49)/625=301/625.